# Symmetry-enriched topological order and quasi-fractonic behavior in $\mathbb{Z}_{N}$ stabilizer codes

### v1

@@ Abstract
We study a broad class of qudit stabilizer codes, termed $\mathbb{Z}_{N}$ bivariate-bicycle (BB) codes, arising either as two-dimensional realizations of modulated gauge theories or as $\mathbb{Z}_{N}$ generalizations of binary BB codes. Our central finding, derived from the polynomial representation, is that the essential topological properties of these $\mathbb{Z}_{N}$ codes can be determined by the properties of their $\mathbb{Z}_{p}$ counterparts, where $p$ are the prime factors of $N$, even when $N$ contains prime powers ($N=\prod_{i}p_{i}^{k_{i}}$). This result yields a significant simplification by leveraging the well-studied framework of codes with prime qudit dimensions. In particular, this insight directly enables the generalization of the algebraic-geometric methods (e.g., the Bernstein-Khovanskii-Kushnirenko theorem) to determine anyon fusion rules in the general qudit situation. Moreover, we analyze the model’s symmetry-enriched topological order (SET) to reveal a quasi-fractonic behavior, resolving the anyon mobility puzzle in this class of models. We also present a computational algebraic method using Gröbner bases over the ring of integers to efficiently calculate the topological order and its SET properties.
@@ I Introduction
Topological order offers a unifying framework for understanding exotic phases of quantum matter indistinguishable by local order parameters. Characterized by long-range entanglement, emergent gauge fields, robust ground state degeneracy, and anyonic excitations with nontrivial braiding statistics [1, 2, 3, 4, 5, 6, 7], it also offers a promising proposals for fault-tolerant quantum computation by encoding quantum information in nonlocal degrees of freedom [8, 9].
The toric code furnishes a simple and illustrative example of a physical system exhibiting topological order and provides a paradigmatic model for fault‑tolerant quantum computation [3]. Recently, there has been an increasing interest in generalizing the toric code in both the condensed matter and the quantum computing communities, with each following distinct yet complementary paths.
@@
From the condensed matter perspective, generalizations of the toric code have been motivated by exploring symmetry structures and novel phases of matter. A natural starting point is to generalize binary systems to $\mathbb{Z}_{N}$ systems describing more general Abelian symmetries. In particular, modulated symmetries are global symmetries dependent on the spatial details of the lattice [52, 53, 54, 55, 56, 57, 58, 59]. Exploring modulated symmetries leads to the discovery of rich phenomena, including Hilbert space fragmentation [60, 61, 62] and UV/IR mixing [63, 64, 65]. A remarkable example of modulated symmetries is the subsystem symmetry, which gives rise to fractons that are immobile in isolation [66, 67, 68, 69, 70, 71, 72, 73, 74, 75]. These insights from modulated symmetries have inspired further generalizations of the toric code [53, 54, 55], uncovering unconventional properties beyond conventional topological order [76, 50]. These generalizations share many similarities with those considered in the quantum computing community. We expect similar methods to be applicable to both studies. However, in contrast to quantum computing, discussions beyond the binary situation present additional challenges.
@@
In this work, we consider various generalizations of toric codes for $\mathbb{Z}_{N}$ qudits exhibiting modulated symmetries. These models fall into a broad class of stabilizer codes we construct, which we refer to as $\mathbb{Z}_{N}$ bivariate-bicycle (BB) codes. We employ the polynomial representation [77] to describe their topological orders. Notably, we find that essential topological order properties of $\mathbb{Z}_{N}$ qudit codes, including topological condition and fusion rules, are immediately determined once we obtain these properties for the $\mathbb{Z}_{p}$ counterparts of this code, where $p$ are the prime factors of $N$. This is a significant simplification, as topological order of stabilizer codes with prime qudit dimensions is widely studied. Specifically, this insight enables us to generalize the algebraic-geometric method of Ref. [50], which uses the Bernstein-Khovanskii-Kushnirenko (BKK) theorem to determine fusion rules, from qubits to the general qudit cases. Moreover, we reveal the quasi-fractonic behavior of the qudit BB code, which is rooted in its symmetry-enriched topological order [78, 79, 80, 81, 82], thereby resolving the mobility puzzle from Ref. [54]. Furthermore, we present a computational algebraic method using the Gröbner bases over the ring of integers to efficiently calculate the topological order and its symmetry-enriched properties of qudit BB codes.
@@
This paper is structured as follows. In Sec. II, we introduce the models considered from a gauging perspective, followed by a review of the polynomial representation of stabilizer codes in Sec. III. We then analyze BB codes, discussing their topological orders in Sec. IV and their quasi-fractonic behaviors in Sec. V. Next, Sec. VI examines the topological orders and quasi-fractonic behaviors of the models introduced in Sec. II. The computational algebraic method used to obtain these results is presented in Sec. VII. Finally, we summarize and discuss our findings in Sec. VIII.
@@ II Models from gauging symmetries
$X=\sum_{j\in\mathbb{Z}_{N}}|j+1\rangle\langle j|,\quad\quad Z=\sum_{j\in\mathbb{Z}_{N}}\omega^{j}|j\rangle\langle j|,$ (1)
where $\omega=e^{\frac{2\pi i}{N}}$. For brevity, we refer to “generalized Pauli” simply as “Pauli” below. They satisfy $X^{N}=Z^{N}=1$ and the commutation relation $ZX=\omega XZ$.
@@
We focus on stabilizer models of the Calderbank-Shor-Steane (CSS) type, each defined by a Hamiltonian of the form
@@ II.1 Watanabe-Cheng-Fuji model
For the case where $a$ is coprime with $N$, the WCF model can be naturally interpreted as a generalized lattice gauge theory, emerging from the process of gauging an exponential symmetry. Explicitly, we follow the procedure described in Ref. [74]. We introduce $\mathbb{Z}_{N}$ qudits (as the matter field) on the vertices of the lattice, governed by the trivial paramagnetic Hamiltonian
@@ II.2 Delfino-Chamon-You model
Another intriguing example related to exponential symmetry is the Delfino-Chamon-You (DCY) model [54], whose excitation mobility still remains a puzzle, and will be addressed in this paper. Its Hamiltonian is given by
@@
where $a\in\mathbb{Z}_{N}$.
Figure 2: Illustration of the DCY model. (a) the $X$-type stabilizer $A_{(i,j)}^{\operatorname{DCY}}$. (b) the $Z$-type stabilizer $B_{(i,j)}^{\operatorname{DCY}}$. In (a) and (b), the location of the stabilizers is marked by the coordinate $(i,j)$ in each figure, while their constituent Pauli operators are shown without explicit coordinates. (c) The minimal coupling terms between gauge and matter qudits. (d) The zero flux condition (yellow) and the gauging symmetry operators (light blue) with all coordinate labels omitted. Throughout the figures, deep blue dots, orange dots, red squares, and purple squares represent the $X$, $Z$, $\sigma^{z}$, and $\sigma^{x}$ operators, respectively.
@@
including the uniform symmetry $U_{0}$, the exponential symmetry $U_{\operatorname{exp}}$ discussed previously, and two additional symmetries with exponential modulation along one spatial coordinate each [83]. Accordingly, symmetric operators are generated by $(\sigma_{(i,j)}^{z})^{a}(\sigma_{(i+1,j)}^{z})^{-(a+1)}\sigma_{(i+2,j)}^{z}$ and $(\sigma_{(i,j)}^{z})^{a}(\sigma_{(i,j+1)}^{z})^{-(a+1)}\sigma_{(i,j+2)}^{z}$, to which gauge field qudits $Z_{(i+\frac{1}{2},j)}$ and $Z_{(i,j+\frac{1}{2})}$ are associated. Applying the gauge procedure results in the DCY model. See Fig. 2(c) and (d).
II.3 Qudit BB codes and product codes
@@
where $c_{i^{\prime},j^{\prime}}$ and $d_{i^{\prime},j^{\prime}}\in\mathbb{Z}_{N}$ specify the powers of Pauli matrices on the lattice edges. We assume only finitely many of the coefficients $c_{i^{\prime},j^{\prime}}$ and $d_{i^{\prime},j^{\prime}}$ are nonzero, ensuring that each stabilizer is locally supported on the lattice and the Hamiltonian remains local. We refer to them as qudit bivariate-bicycle (BB) codes, which generalize the qubit BB codes introduced in Ref. [26] to the $\mathbb{Z}_{N}$ setting.
If $c_{i^{\prime},j^{\prime}}=0$ for $j^{\prime}=0$ and $d_{i^{\prime},j^{\prime}}=0$ for $i^{\prime}=0$, we refer to this type of BB code as product codes, since they can be constructed from the tensor product of two classical codes from the quantum computing perspective. Explicitly, the stabilizers of product codes take the form
$A_{(i,j)}$ $=\prod_{i^{\prime},j^{\prime}}X^{c_{0,j^{\prime}}}_{(i-i^{\prime}+\frac{1}{2},j)}X^{d_{i^{\prime},0}}_{(i,j-j^{\prime}+\frac{1}{2})},$ (13)
$B_{(i,j)}$ $=\prod_{i^{\prime},j^{\prime}}Z^{d_{i^{\prime},0}}_{(i+i^{\prime}+\frac{1}{2},j)}Z^{-c_{0,j^{\prime}}}_{(i,j+j^{\prime}+\frac{1}{2})},$
@@ III Polynomial representation
The polynomial ring formalism is set up through the Laurent polynomial ring $R=\mathbb{Z}_{N}[x^{\pm},y^{\pm}]$, which can be viewed as the group ring $\mathbb{Z}_{N}[\Lambda]$. Here, $\Lambda=\{x^{i}y^{j}|i,j\in\mathbb{Z}\}$ denotes the lattice translation group, with $(i,j)$ locating unit cells. Each unit cell of BB codes contains two qudits, located at $(i+\frac{1}{2},j)$ and $(i,j+\frac{1}{2})$. Within this framework, Pauli operators are represented as vectors in $R^{4}$, in particular,
@@
For two multi-qudit Pauli operators represented as $v_{1},v_{2}\in R^{4}$, they commute with each other if and only if the constant term of the symplectic bilinear form ${v}_{1}^{\dagger}\lambda_{q}v_{2}$ vanishes. Here, $v_{1}^{\dagger}=\bar{v}_{1}^{T}$ where $T$ denotes the transpose operation and $\bar{v}_{1}(x,y)=v_{1}(\bar{x},\bar{y})$ with $\bar{x}=x^{-1}$ and $\bar{y}=y^{-1}$, and $\lambda_{q}$ denotes the standard symplectic matrix, given by
@@
In particular, the first and second columns of $\partial$ can generate all $X$ type stabilizers $A_{(i,j)}$ and $Z$-type stabilizers $B_{(i,j)}$ by translation actions, respectively. The symplectic bilinear form of the first and second columns of $\partial$ equals 0, so all stabilizers of BB codes commute with each other, and BB codes are well defined.
@@
All stabilizers commute with the Hamiltonian, ensuring that $\partial\circ\varepsilon=0$. Formally, this means that the stabilizer map $\partial$ and the excitation map $\varepsilon$ form the following chain complex:
@@ IV.1 Topological condition
In other words, the chain complex Eq. (18) is exact. Given the explicit form of $\partial$ and $\varepsilon$ in Eq. (16) and Eq. (17), respectively, the $\mathbf{BB}_{N}(f,g)$ code is topological if and only if $\ker\varepsilon_{A}=\operatorname{im}\partial_{B}$, or equivalently, $\ker\varepsilon_{B}=\operatorname{im}\partial_{A}$.
@@
We present the proof in Appendix A. Applying Theorem 1 leads to a convenient criterion for topological order, presented in the following corollary.
@@
$\mathcal{Q}=\frac{R}{(f,g)},$ (22)
is finite.
The detailed proof of this criterion is offered in Appendix A. As we discuss in the following section, if the BB code is topological, $\mathcal{Q}$ describes the anyon types of electric charges of this model, implying the total number of anyon types is finite. This result aligns with the general result that all topological excitations of translationally invariant 2D stabilizer codes are mobile [84].
@@ IV.2 Anyon types and fusion Rules
since $\operatorname{im}\varepsilon$ captures precisely those excitation patterns that are realizable by local operators, and two elements $\omega,\omega^{\prime}\in R^{2}$ represent two excitation patterns that can transform into each other by local operators if and only if they are equivalent in $\mathcal{C}$. The quotient modules $\mathcal{Q}$ and $\overline{\mathcal{Q}}$ in Eq. (23) specify the $X$ and $Z$-type excitations of BB codes. Physically, they correspond the electric and magnetic charges of the model are in the emergent gauge theory, respectively.
@@
In what follows, we refer to $Q_{i}$ as the topological index. The proof of Theorem 2 involves basic homological algebra and is detailed in Appendix C. Theorem 2 provides a convenient and efficient way to evaluate the anyon types and their fusion rules of BB codes using results from prime qudit dimensions. Specifically, this enables us to extend the algebraic-geometric approach for determining fusion rules based on the Bernstein-Khovanskii-Kushnirenko (BKK) theorem. This method was previously limited to qubit BB codes [50]; it can now be applied to composite qudit dimensions. See Sec. VI.3 and Sec. VI.4 for two concrete examples of applications of Theorem 2.
V Quasi-Fractonic behavior and SET orders of BB Codes
In this section, we study the mobility of anyons in BB codes. We examine the quasi-fractonic behavior of BB codes and how these behaviors reflect the system’s symmetry-enriched topological (SET) order. Additionally, we discuss the ground state degeneracy of BB codes, whose size dependence property serves as a key feature of the translation SET.
V.1 Quasi-fractonic behavior of BB codes

### v2

# Symmetry-enriched topological order and quasifractonic behavior in $\mathbb{Z}_{N}$ stabilizer codes

@@ Abstract
We study a broad class of qudit stabilizer codes, termed $\mathbb{Z}_{N}$ bivariate-bicycle (BB) codes, arising either as two-dimensional realizations of modulated gauge theories or as $\mathbb{Z}_{N}$ generalizations of binary BB codes. Our central finding, derived from the polynomial representation, is that the essential topological properties of these $\mathbb{Z}_{N}$ codes can be determined by the properties of their $\mathbb{Z}_{p}$ counterparts, where $p$ are the prime factors of $N$, even when $N$ contains prime powers ($N=\prod_{i}p_{i}^{k_{i}}$). This result yields a significant simplification by leveraging the well-studied framework of codes with prime qudit dimensions. In particular, this insight directly enables the generalization of the algebraic-geometric methods (e.g., the Bernstein-Khovanskii-Kushnirenko theorem) to determine anyon fusion rules in the general qudit situation. Moreover, we elucidate the symmetry-enriched topological (SET) order underlying the quasifractonic behavior in qudit BB codes (including the Delfino-Chamon-You model), resolving the associated anyon mobility puzzle. We also develop an efficient computational algebraic method, based on Gröbner bases over the ring of integers, to determine both the topological order and its SET properties.
@@ I Introduction
Topological order offers a unifying framework for understanding exotic phases of quantum matter indistinguishable by local order parameters. Characterized by long-range entanglement, emergent gauge fields, robust ground state degeneracy, and anyonic excitations with nontrivial braiding statistics [1, 2, 3, 4, 5, 6, 7], it also offers promising proposals for fault-tolerant quantum computation by encoding quantum information in nonlocal degrees of freedom [8, 9].
The toric code furnishes a simple and illustrative example of a physical system exhibiting topological order and provides a paradigmatic model for fault-tolerant quantum computation [3]. Recently, there has been increasing interest in generalizing the toric code in both the condensed matter and the quantum computing communities, with each following distinct yet complementary paths.
@@
From the condensed matter perspective, generalizations of the toric code have been motivated by exploring symmetry structures and alternative phases of matter. A natural starting point is to generalize binary systems to $\mathbb{Z}_{N}$ systems describing more general Abelian symmetries. In particular, modulated symmetries are global symmetries dependent on the spatial details of the lattice [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. Exploring modulated symmetries leads to the discovery of rich phenomena, including Hilbert space fragmentation [64, 65, 66] and UV/IR mixing [67, 68, 69, 70, 71, 72, 73]. A remarkable example of modulated symmetries is the subsystem symmetry, which gives rise to fractons that are immobile in isolation [74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92]. These insights from modulated symmetries have inspired further generalizations of the toric code [53, 54, 55], uncovering unconventional properties beyond conventional topological order [93, 50]. These generalizations share many similarities with those considered in the quantum computing community. We expect similar methods to be applicable to both studies. However, in contrast to quantum computing, discussions beyond the binary situation present additional challenges.
@@
In this work, we consider various generalizations of toric codes for $\mathbb{Z}_{N}$ qudits exhibiting modulated symmetries. These models fall into a broad class of stabilizer codes we construct, which we refer to as $\mathbb{Z}_{N}$ BB codes. We employ the polynomial representation [94] to describe their topological orders. Notably, we find that the essential topological-order properties of $\mathbb{Z}_{N}$ qudit codes—including the topological condition and the fusion rules—are determined once these properties are known for the corresponding $\mathbb{Z}_{p}$ codes, where $p$ runs over the prime factors of $N$. This is a significant simplification, as the topological orders of stabilizer codes with prime qudit dimensions have been well studied. Specifically, it enables us to generalize the algebraic-geometric method of Ref. [50], which uses the Bernstein-Khovanskii-Kushnirenko (BKK) theorem to determine fusion rules, from qubits to the general qudit cases. In addition, we elucidate the symmetry-enriched topological (SET) structure [95, 96, 97, 98, 99] underlying the quasifractonic behavior. We develop an efficient computational algebraic method, based on the Gröbner bases over integers, to determine both the topological order and its symmetry-enriched properties for qudit BB codes. As an application, our framework provides a precise characterization of the Delfino-Chamon-You (DCY) model [54] and resolves its mobility puzzle regarding whether one-to-one anyon hopping is permitted over any finite distance.
@@
This paper is structured as follows. In Sec. II, we introduce the models considered from a gauging perspective, followed by a review of the polynomial representation of stabilizer codes in Sec. III. We then analyze BB codes, discussing their topological orders in Sec . IV and their quasifractonic behaviors in Sec. V. Next, Sec. VI examines the topological orders and quasifractonic behaviors of the models introduced in Sec. II. The computational algebraic method used to obtain these results is presented in Sec. VII. Finally, we summarize and discuss our findings in Sec. VIII.
@@ II Models from gauging symmetries
$X=\sum_{j\in\mathbb{Z}_{N}}|j\rangle\langle j+1|,\quad\quad Z=\sum_{j\in\mathbb{Z}_{N}}\omega^{j}|j\rangle\langle j|,$ (1)
where $\omega=e^{\frac{2\pi i}{N}}$. For brevity, we refer to generalized Pauli simply as Pauli below. They satisfy $X^{N}=Z^{N}=1$ and the commutation relation $XZ=\omega ZX$.
@@
We focus on stabilizer models of the Calderbank-Shor-Steane type, each defined by a Hamiltonian of the form
@@ II.1 Watanabe-Cheng-Fuji model
For the case where $a$ is coprime to $N$, the WCF model can be naturally interpreted as a generalized lattice gauge theory, emerging from the process of gauging an exponential symmetry. Explicitly, we follow the procedure described in Ref. [83]. We introduce $\mathbb{Z}_{N}$ qudits (as the matter field) on the vertices of the lattice, governed by the trivial paramagnetic Hamiltonian
@@ II.2 Delfino-Chamon-You model
Another intriguing example related to exponential symmetry is the DCY model [54], whose excitation mobility still remains a puzzle, and will be addressed in this paper. Its Hamiltonian is given by
@@
where $a\in\mathbb{Z}_{N}$. Note that there are two qudits per unit cell, which we take to be located at $(i-\frac{1}{2},j)$ and $(i,j-\frac{1}{2})$; by shifting them to the lattice vertices, the model may be transformed into precisely the form presented in Ref. [54].
Figure 2: Illustration of the DCY model. (a) $X$-type stabilizer $A_{(i,j)}^{\operatorname{DCY}}$. (b) $Z$-type stabilizer $B_{(i,j)}^{\operatorname{DCY}}$. In (a) and (b), the location of the stabilizers is marked by the coordinate $(i,j)$ in each figure, while their constituent Pauli operators are shown without explicit coordinates. (c) The minimal coupling terms between gauge and matter qudits. (d) The zero flux condition (yellow) and the gauging symmetry operators (light blue) with all coordinate labels omitted. Throughout the figures, deep blue dots, orange dots, red squares, and purple squares represent the $X$, $Z$, $\sigma^{z}$, and $\sigma^{x}$ operators, respectively.
@@
including the uniform symmetry $U_{0}$, the exponential symmetry $U_{\operatorname{exp}}$ discussed previously, and two additional symmetries with exponential modulation along one spatial coordinate each [100]. If these global symmetries are independent, symmetric operators are generated by $(\sigma_{(i,j)}^{z})^{a}(\sigma_{(i+1,j)}^{z})^{-(a+1)}\sigma_{(i+2,j)}^{z}$ and $(\sigma_{(i,j)}^{z})^{a}(\sigma_{(i,j+1)}^{z})^{-(a+1)}\sigma_{(i,j+2)}^{z}$, to which gauge field qudits $Z_{(i+\frac{1}{2},j)}$ and $Z_{(i,j+\frac{1}{2})}$ are associated. Applying the gauge procedure results in the DCY model. See Fig. 2(c) and (d).
II.3 Qudit BB codes and hypergraph product codes
@@
where $c_{i^{\prime},j^{\prime}}$ and $d_{i^{\prime},j^{\prime}}\in\mathbb{Z}_{N}$ specify the powers of Pauli matrices on the lattice edges. We assume only finitely many of the coefficients $c_{i^{\prime},j^{\prime}}$ and $d_{i^{\prime},j^{\prime}}$ are nonzero, ensuring that each stabilizer is locally supported on the lattice. We refer to them as qudit BB codes, which generalize the qubit BB codes introduced in Ref. [26] to the $\mathbb{Z}_{N}$ setting.
If $c_{i^{\prime},j^{\prime}}=0$ for $j^{\prime}\neq 0$ and $d_{i^{\prime},j^{\prime}}=0$ for $i^{\prime}\neq 0$, we refer to this type of BB code as a hypergraph product (HGP) code, since it arises from two classical codes via a tensor product construction in quantum error correction [14]. Explicitly, the stabilizers of HGP codes take the form
$A_{(i,j)}$ $=\prod_{i^{\prime},j^{\prime}}X^{c_{i^{\prime},0}}_{(i-i^{\prime}+\frac{1}{2},j)}X^{d_{0,j^{\prime}}}_{(i,j-j^{\prime}+\frac{1}{2})},$ (13)
$B_{(i,j)}$ $=\prod_{i^{\prime},j^{\prime}}Z^{d_{0,j^{\prime}}}_{(i+i^{\prime}+\frac{1}{2},j)}Z^{-c_{i^{\prime},0}}_{(i,j+j^{\prime}+\frac{1}{2})},$
@@ III Polynomial representation
The polynomial ring formalism is set up through the Laurent polynomial ring $R=\mathbb{Z}_{N}[x^{\pm},y^{\pm}]$, which can be viewed as the group ring $\mathbb{Z}_{N}[\Lambda]$. Here, $\Lambda=\{x^{i}y^{j}|i,j\in\mathbb{Z}\}$ denotes the lattice translation group, with $(i,j)$ locating unit cells. Here, we choose the convention that the unit cell contains two qudits at $(i+\frac{1}{2},j)$ and $(i,j+\frac{1}{2})$. Within this framework, Pauli operators are represented as vectors in $R^{4}$, in particular,
@@
For two multi-qudit Pauli operators represented as $v_{1},v_{2}\in R^{4}$, they commute with each other if and only if the constant term of the symplectic bilinear form ${v}_{1}^{\dagger}\lambda_{q}v_{2}$ vanishes. Here, $v_{1}^{\dagger}=\bar{v}_{1}^{T}$ where $T$ denotes the transpose operation and $\bar{v}_{1}(x,y)=v_{1}(\bar{x},\bar{y})$ with $\bar{x}=x^{-1}$ and $\bar{y}=y^{-1}$, and $\lambda_{q}$ denotes the standard symplectic form matrix, given by
@@
In particular, the first and second columns of $\partial$ can generate all $X$-type stabilizers $A_{(i,j)}$ and $Z$-type stabilizers $B_{(i,j)}$ by translation actions, respectively. The symplectic bilinear form of the first and second columns of $\partial$ equals 0, so all stabilizers of BB codes commute with each other, and BB codes are well defined.
@@
All stabilizers commute with the Hamiltonian, ensuring that $\varepsilon\circ\partial=0$. Formally, this means that the stabilizer map $\partial$ and the excitation map $\varepsilon$ form the following chain complex:
@@ IV.1 Topological condition
In other words, the chain complex Eq. (18) is exact. Given the explicit form of $\partial$ and $\varepsilon$ in Eqs. (16) and (17), respectively, the $\mathbf{BB}_{N}(f,g)$ code is topological if and only if $\ker\varepsilon_{A}=\operatorname{im}\partial_{B}$, or equivalently, $\ker\varepsilon_{B}=\operatorname{im}\partial_{A}$.
@@
We present the detailed proof in Appendix A. Applying Theorem 1 leads to a convenient criterion for topological order, presented in the following corollary.
@@
$\mathcal{Q}=\frac{R}{(f,g)}$ (22)
is finite, where $(f,g)$ denotes the ideal generated by polynomials $f$ and $g$.
The proof of this criterion is offered in Appendix A. As we discuss in the following section, if the BB code is topological, $\mathcal{Q}$ describes the anyon types of the electric charge sector of this model, implying the total number of anyon types is finite. This result aligns with the general mobility theorem for topological excitations of translationally invariant 2D stabilizer codes in Ref. [101].
@@ IV.2 Anyon types and fusion rules
since $\operatorname{im}\varepsilon$ captures precisely those excitation patterns that are realizable by local operators, and two elements $\omega,\omega^{\prime}\in R^{2}$ represent two excitation patterns that can transform into each other by local operators if and only if they are equivalent in $\mathcal{C}$. The sectors $\mathcal{Q}$ and $\overline{\mathcal{Q}}$ in Eq. (23) specify the $X$ and $Z$-type excitations of BB codes. Physically, they correspond to the electric and magnetic charges of the model in the emergent gauge theory, respectively.
@@
In what follows, we refer to $Q_{i}$ as the topological index. The proof of Theorem 2 involves basic homological algebra and is detailed in Appendix C. Theorem 2 provides a convenient and efficient way to evaluate the anyon types and their fusion rules of BB codes using results from prime qudit dimensions. Specifically, this enables us to extend the algebraic-geometric approach for determining fusion rules based on the BKK theorem. This method was previously limited to qubit BB codes [50]; it can now be applied to composite qudit dimensions. See Secs. VI.3 and VI.4 for two concrete applications of Theorem 2.
V Quasifractonic behavior and SET orders of BB Codes
In this section, we analyze anyon hopping and the GSD, revealing how their distinctive features emerge from the interplay between topological order and translation symmetry.
V.1 Quasifractonic mobility of anyons